# Learning Object Compliance via Young’s Modulus from Single Grasps using Camera-Based Tactile Sensors

### v2

@@ Abstract
Compliance is a useful parametrization of tactile information that humans often utilize in manipulation tasks. It can be used to inform low-level contact-rich actions or characterize objects at a high-level. In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across object shape and material. Using camera-based tactile sensors, proprioception, and force measurements, we present a novel approach to estimate object compliance as Young’s modulus $E$ from parallel grasps. We evaluate our method over a novel dataset of 285 common objects, which includes a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa. Data is collected over automated parallel grasps of each object. Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping. This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset. This is a drastic improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively. Importantly, this estimation system performs irrespective of object geometry and demonstrates robustness across object materials. Thus, it could be applied in a general robotic manipulation setting to characterize unknown objects and inform decision-making, for instance to sort produce by ripeness.
Keywords: Robotic Manipulation, Tactile Sensing, Compliance
1 Introduction
General and robust object compliance estimation could significantly enhance robotic perception systems. For instance, a ripe avocado appears visually identical to an unripe avocado, yet their compliance differs greatly. Without the ability to accurately estimate compliance, a robot cannot effectively distinguish between such visually similar objects. Compliance measurement has proven valuable in multiple applications, including identifying the ripeness of fruits and vegetables [1, 2, 3, 4] and in biomedical fields [5, 6, 7]. In the future, a universal compliance estimation architecture could be applied in assembly or recycling processes to inform decisions based on object material.
In robotics, we seek to replicate the human ability to perceive properties like compliance through touch [8]. Humans naturally use tactile information to characterize objects and inform interactions [9]. In the case of compliance, we sense discrepancies between soft and hard objects through fingertip pressure and deformation [10]. High-resolution, camera-based tactile sensors like GelSight are specifically designed to track the deformation of a robotic fingertip during contact [11, 12].
However, even with the help of high-resolution tactile sensing, achieving human-like compliance perception remains challenging, especially in unstructured environments. An object’s deformation in response to loading depends on both its shape and material composition. Thus, without known geometry and pose tracking, analytical methods cannot be utilized to rigorously determine compliance [13]. Alternatively, learning-based implementations have been developed to estimate compliance from tactile images alone without the use of mechanical modeling [14, 5, 15, 1]. While these methods have demonstrated relative success in estimating compliance independent of shape, they still face limitations in describing compliance across a wide range of materials.
We aim to create a compliance estimation system that is robust to both shape and material by integrating analytical modeling with machine learning. Young’s modulus ($E$) is a fundamental property in mechanical modeling that offers a measure of compliance for nearly all materials [16]. By parameterizing compliance using Young’s modulus, we can seamlessly incorporate analytical models and create estimates of compliance on a universal scale. Our hybrid approach leverages the strengths of analytical modeling to better generalize across materials, while using learning techniques to address complexities posed by diverse geometries. Specifically, we use a neural network to fuse CNN-encoded tactile images with estimates from analytical contact models. We train our approach across a novel dataset of 285 common objects, including a wide variety of shapes and materials.
Our primary contribution is a novel compliance estimation archiecture which demonstrates increased accuracy over the current state-of-the-art. Additionally, we contribute a novel dataset which could be used in the development of future compliance estimation architectures. Using our diverse dataset, we validate that our hybrid estimation approach can generalize well across object shape and material. Thus, our estimation system could be applied in general robotic manipulation to characterize unknown objects and inform decision-making. This is especially beneficial in scenarios where visually similar objects may have distinct material composition.
@@ 1 Introduction
2 Related Work
2.1 GelSight Sensors & Contact Modeling
@@
Camera-based tactile sensors, such as GelSight sensors, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [12, 17, 18, 19, 20, 21]. They are highly effective for sensing texture [22, 23, 24], contact force [12], friction [25], and object shape and pose [26, 27]. However, reliably modeling contact with these soft sensors remains challenging. Simulators have been developed to generate tactile images and deformed geometry of sensors using pre-computed finite element method (FEM) solutions [28]. Finite element analysis has proven a viable way to model soft contact on hardware and used to estimate forces from tactile data [29, 30]. Though modeling with FEM is accurate, it requires a precise map of object geometry and solutions often must be pre-computed offline.
@@
Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [31, 32]. This closed-form method offers improved computational efficiency over FEM and simulated methods. Still, it requires assumptions on contact geometry, namely that geometry should be axisymmetric. Hertzian contact models have been utilized for force estimation [33, 5] and surface curvature detection from tactile data [34]. These closed-form constitutive models can easily be inverted to determine object compliance.
2.2 Compliance Estimation
@@
Estimating the compliance of a contacted object remains a complex problem in robotic manipulation, particularly without assumptions on object shape or material type. Analytical approaches to compliance estimation often lack robustness to variations in geometry. A traditional approach to model compliance is to apply Hooke’s Law using measurements of grasping force and positions [13]. Similar models have been foundational in developing low-level grasping control policies [35, 36]. This method can provide an effective stiffness of grasping, but may struggle when applied to unknown objects with complex shapes. By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [37, 38, 39]. These methods have a stronger ability to sense and generalize across contact geometry but still struggle to account for complex contact geometries.
@@
Recent learning-based approaches to compliance estimation have demonstrated the ability to better generalize across geometry but struggle when applied over a wide-range of material types. Data-driven architectures have been developed to estimate the Shore 00 hardness of an object in contact with a GelSight sensor directly from tactile images without any mechanical modeling [40, 14]. The Shore 00 hardness scale is designed to quantify the compliance of soft rubbers [41]. This model is trained using exclusively rubber objects and is thus fundamentally limited to this material domain. Alternatively, LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of ‘soft’ or ‘hard’ [15, 1]. These large data models are promising but have not been used to assess finer differences in compliance on a continuous scale. Other data-driven implementations have been developed using tactile sensing to directly measure the ripeness of produce [2, 3, 4]. As before, these methods are restricted by their dataset and labeling method. In all cases, labeling choices inhibit models from integrating analytical modeling alongside learning.
3 Methods
3.1 Overview
We develop a hybrid system that fuses analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp. Through analytical modeling, we can create a well-founded, preliminary estimate for Young’s modulus. By incorporating learning, we can compensate for assumptions taken by analytical models to better generalize across contact geometry. A large, diverse dataset of 285 common objects is collected to train and evaluate our method. Our approach of combining analytical models with data-driven methods is partially inspired by residual physics architectures [42] and physics-informed tactile models [43].
Our hybrid estimation architecture is displayed in Fig. 2. After grasping, tactile images and measurements of force $F$ and gripper width $w$ are fed into analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$. In parallel, a selected number $N_{\text{frames}}$ of tactile images from a single sensor are processed through a CNN. Features extracted from these images are concatenated with grasp measurements $F$, $w$ into a large fully-connected decoder. This multi-tower network is similar to other grasp-based architectures [44]. Output features from the large decoder are sent to a smaller decoder alongside analytical estimates to produce a final hybrid estimate of Young’s modulus $\hat{E}$.
@@ 3.1 Overview
Figure 2: Young’s modulus estimation architecture. Tactile images, depth, and measured grasping force $F$ and width $w$ are sent into analytical and learned models. Estimates from analytical models are fed into a fully-connected decoder alongside learned grasp features.
@@
Code for the project is publicly available on GitHub. Collected data for multiple grasps of all objects with labeled Young’s moduli and Shore hardness is publicly available on HuggingFace.
3.2 Analytical Contact Models
@@
Simple Elasticity: Parallel grasping can be modeled as uniaxial loading using an adaptation of Hooke’s Law. A diagram of parallel grasping is provided in Fig. 3. At every instance $t$, force $F$ is applied to the grasped object, with gripper at width $w$. Traditional methods have used these measurements alone to estimate the compliance or stiffness of grasped objects [13].
@@
With camera-based tactile sensors, we can increase the fidelity of the model by measuring contact area $A$ and depth $d$ [12]. Contact area $A$ is determined by masking depth images with constant threshold of 0.1mm. This threshold was chosen empirically to cancel out potential noise. Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume constant contact area ($A(t)=A_{\text{sensor}}$). We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq. 1. This model assumes symmetry between parallel fingers. We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than an empirically-defined threshold.
$\sigma(t)=\frac{F(t)}{A(t)}\;,\quad\quad\epsilon(t)=\frac{\Delta w(t)+2\Delta d(t)}{w(0)+2d(0)}=\frac{(w(t)+2d(t))-(w(0)+2d(0))}{w(0)+2d(0)}$ (1)
@@
In the elastic regime, stress $\sigma$ and $\epsilon$ are related by Young’s modulus [16]. Using measurements across grasping time $t$, we fit an estimate for Young’s modulus $\hat{E}_{\text{elastic}}$ following Eq. 2.
$\boldsymbol{\sigma}=\hat{E}_{\text{elastic}}\boldsymbol{\epsilon}$ (2)
@@
Despite its simplicity, this method does not properly consider object geometry and implicitly assumes a consistent profile along the grasping axis. Furthermore, it overlooks precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
@@
Figure 3: Simple mechanics of parallel grasping. Spherical object of radius $R$ is grasped with soft GelSight sensors. Normal force $F$ is applied at width $w$. Penetration depth $d$ and contact radius $a$ can be measured with tactile sensors. Timestep $t=0$ is the moment of first contact over the grasp.
@@
Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact [45]. In a simple case, contact is assumed to be without friction or adhesion [46, 47]. Unlike the previous method, Hertzian mechanics closely analyzes geometry through contact. It has been applied to model soft contact in robotics [18, 5], including for GelSight-like tactile sensors [31, 34, 18, 5].
@@
GelSight sensors are made with highly elastic silicone rubber [17]. Unless a grasped object is fully rigid, we expect both bodies to deform through contact. This implies the maximum measured penetration depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$. This phenomenon is depicted in Fig. 3. An aggregate modulus $E^{*}$ describes the bodies’ combined resistance to deformation in Eq. 3, which depends on the Poisson’s ratio $\nu$ of each body.
@@
Hertzian solutions have been derived for a wide range of contact geometries [48]. We model the sensor as a flat plane. Although we are analyzing contact with an object whose shape is unknown, we model the geometry of all grasped objects as spherical. This simplification is reasonable for convex objects, but may produce bounded errors over our geometrically diverse dataset. Our neural network can learn to compensate for this intrinsic and repeatable geometric modeling error.
Since we assume contact profiles are spherical, we calculate contact radius $a$ based on observed total contact area $A$. Additionally, we calculate the apparent radius of the unknown object $R$ using contact area and gripper displacement [48]. Using these parameters, we can relate depth $d$ to normal force $F$ via the method of dimensionality reduction (MDR) [48, 49], as formulated in Eq. 4. More detailed derivation for the MDR is provided in Appendix A. With these equations established and measurements over the duration of a grasp, we apply least-squares to retrieve a best fit value $\hat{E}^{*}$.
$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)},\quad\boldsymbol{d}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{{\hat{E}}^{*2}}\boldsymbol{F}}{32\boldsymbol{R}^{2}}\right)}^{\frac{1}{3}}\boldsymbol{a}$ (4)
We compute an estimated modulus for our unknown object $\hat{E}_{\text{hertz}}$ through Eq. 5, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$. We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [50]. After analytical estimates are computed, linear scaling is added to account for potential constant errors across the dataset.

### v3

@@ Abstract
Compliance is a useful parametrization of tactile information that humans often utilize in manipulation tasks. It can be used to inform low-level contact-rich actions or characterize objects at a high-level. In robotic manipulation, existing approaches to estimate compliance have struggled to generalize across both object shape and material. Using camera-based tactile sensors, proprioception, and force measurements, we present a novel approach to estimate object compliance as Young’s modulus $E$ from parallel grasps. We evaluate our method over a novel dataset of 285 common objects, including a wide array of shapes and materials with Young’s moduli ranging from 5.0 kPa to 250 GPa. Combining analytical and data-driven approaches, we develop a hybrid system using a multi-tower neural network to analyze a sequence of tactile images from grasping. This system is shown to estimate the Young’s modulus of unseen objects within an order of magnitude at 74.2% accuracy across our dataset. This is an improvement over purely analytical and data-driven baselines which exhibit 28.9% and 65.0% accuracy respectively. Importantly, this estimation system performs irrespective of object geometry and demonstrates increased robustness across material types.
I Introduction
Humans naturally use tactile feeedback to characterize objects and inform our interactions with them [1]. Compliance, or softness, is useful in everyday tasks. For example, we can determine the ripeness of an avocado or other fruits by assessing their compliance. Crucially, surgeons rely on tactile feedback and perceived compliance to accurately characterize tissues and detect tumors [2].
In robotics, if we would like to better replicate the natural proficiency of human manipulation through teleoperated action, we must communicate this tactile information to remote users. Recently, haptic devices have been developed to display softness for remote users [3, 4, 5]. However, these displays are only useful if we can reliably provide them a measure of compliance in real-time.
@@ I Introduction
Humans sense discrepancies between soft and hard objects through fingertip pressure and deformation [6]. Camera-based tactile sensors, like GelSight, are specifically designed to track these parameters on robotic fingertips [7, 8]. Analytical models, such as Hertzian mechanics, follow a similar approach to approximate compliance as humans may [9, 10]. Ideally, we could input tactile data to these models to determine compliance. However, traditional analytical frameworks often struggle to generalize across arbitrary geometries that may be encountered in unstructured environments. Since an object’s deformation under load depends on both its shape and compliance, analytical methods cannot rigorously determine compliance without prior knowledge of object geometry [11].
Alternatively, learning-based implementations have been developed to estimate compliance from tactile images alone without the use of mechanical modeling [12, 13, 14, 15]. While these methods have demonstrated success in estimating compliance independent of shape, they still face limitations in describing compliance across a wide range of materials. Many of these models elect to exclusively label the compliance of rubber objects via the Shore hardness scale. To successfully incorporate compliance feedback into a teleoperation interface, we need labels of compliance on a universal scale to replicate the continuity that humans experience in real-world manipulation.
To address these limitations and move towards more generalized compliance estimation, we propose a hybrid approach that combines analytical modeling with learning-based methods to estimate an object’s compliance as Young’s modulus $E$ from a single parallel grasp. By incorporating analytical models, we improve generalization across materials, while using a learned CNN to account for geometric complexities. We choose to label compliance as Young’s modulus to align with analytical models and so that our system produces estimations on a universal scale.
In this work, we present the following contributions:
  • A novel compliance estimation architecture which demonstrates increased accuracy over the current state-of-the-art. Unlike previous methods, our hybrid learned architecture incorporates analytical modeling.
  • Introduce Young’s modulus $E$ as a label of compliance to enable seamless incorporation of analytical models and produce estimates on a universal scale.
  • A novel dataset containing 285 common objects, featuring a wide variety of shapes and materials, which could be used in developing future compliance estimation architectures.
We validate that our hybrid estimation approach can generalize reasonably well across our diverse dataset. Thus, our estimation system could be applied in robotic manipulation systems to more robustly estimate compliance.
II Related Works
II-A GelSight Sensors & Contact Modeling
@@
Camera-based tactile sensors, such as GelSight, are designed to track contact geometry in high-resolution using a camera placed behind an elastomer gel pad [8, 16, 17]. They are effective for sensing texture [18], object shape and pose [19], contact force [8], and friction [20]. However, reliably modeling contact with these soft sensors remains challenging. Finite element method (FEM) has proven a viable option [21, 22]. Although modeling with FEM is accurate, it requires a precise map of object geometry and solutions often must be pre-computed offline.
@@
Alternative approaches have been developed using Hertzian mechanics to model soft contact in robotic grasping [23]. This closed-form method offers improved computational efficiency over FEM and other methods. Still, it requires assumptions on contact geometry, namely that geometry should be axisymmetric. Hertzian contact models have been utilized for force estimation [24, 13] and curvature detection from tactile data [25]. These closed-form constitutive models can be inverted to determine compliance.
II-B Compliance Estimation
@@
Estimating the compliance of a contacted object remains a difficult problem in robotic manipulation, particularly without assumptions on object shape or material type. Analytical approaches to compliance estimation often lack robustness to variations in geometry. The most traditional approach to model compliance is to apply Hooke’s Law to measurements of grasping force and position [11]. This method can provide an effective stiffness of grasping, but may struggle when applied to unknown objects with complex shapes. By incorporating tactile sensing, it is possible to create higher fidelity compliance estimation models [26, 27]. These methods have a stronger ability to generalize across contact geometry through sensing, but still struggle to account for complex contact geometries.
@@
Recent learning-based approaches to compliance estimation have demonstrated the ability to better generalize across geometry, but struggle when applied over a wide-range of material types. Data-driven architectures have been developed to estimate the Shore 00 hardness of an object in contact with a GelSight sensor directly from tactile images without any mechanical modeling [12]. The Shore 00 hardness scale is designed to quantify the compliance of soft rubbers [28]. This model is trained using exclusively rubber objects and is thus limited to this material domain. Alternatively, LLM-backed methods have been developed to classify tactile images with binary semantic descriptions of “soft” or “hard” [14, 15]. These large data models are promising, but have not been used to assess finer differences in compliance on a continuous or quantitative scale. Other data-driven implementations have been developed using tactile sensing to directly detect the ripeness of produce [29, 30, 31]. As before, these methods are restricted by their dataset and labeling method. In all cases, models do not incorporate analytical modeling alongside learning.
@@
Figure 2: Young’s modulus estimation architecture. Tactile images, depth, and measured grasping force $F$ and width $w$ are sent into analytical and learned models. Estimates from analytical models are fed into a fully-connected decoder alongside learned grasp features to produce hybrid estimate $\hat{E}$.
III Methods
III-A Overview
We develop a hybrid system that fuses analytical models alongside learning to estimate the Young’s modulus of general unknown objects from a single parallel grasp. Through analytical modeling, we can create a well-founded, preliminary estimate for Young’s modulus. By incorporating learning, we can compensate for assumptions taken by analytical models to better generalize across contact geometry. A large, diverse dataset of 285 common objects is collected to train and evaluate our method. This approach of combining analytical models with data-driven methods is partially inspired by residual physics architectures [32] and physics-informed tactile models [33].
Our hybrid estimation architecture is shown in Fig. 2. After grasping, tactile images and measurements of normal force $F$ and gripper width $w$ are input into analytical models to fit respective estimates of Young’s modulus $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$. In parallel, a selected number $N_{\text{frames}}$ of tactile images from a single sensor are passed to a CNN. Our CNN architecture is adapted from [34], which has demonstrated strong performance in classifying sequences of tactile images. Features extracted from tactile images by this CNN are concatenated with grasp measurements $F$, $w$ and fed into a large fully-connected decoder. Output features from the large decoder are sent alongside analytical estimates to a smaller decoder to produce a final hybrid estimate of Young’s modulus $\hat{E}$. This multi-tower network is inspired by other grasp-based architectures [35].
@@ III-A Overview
Code for the project is publicly available on GitHub. Collected data with labeled Young’s moduli and Shore hardnesses is publicly available on HuggingFace.
III-B Analytical Contact Models
@@
Simple Elasticity: Parallel grasping can be modeled as uniaxial loading following Hooke’s Law. A diagram of parallel grasping is provided in Fig. 3a. At every instance $t$, normal force $F$ is applied to the grasped object, with gripper at width $w$. Traditional methods have used these measurements alone to estimate the stiffness of grasped objects [11]. Throughout this paper, we assume objects to be made of isotropic materials.
@@
With camera-based tactile sensors, we can increase the fidelity of contact models by measuring contact area $A$ and depth $d$ [8]. Contact area $A$ is determined by masking tactile depth images with a constant threshold of 0.1mm. This threshold was chosen empirically to cancel out potential noise. Without tactile sensing, we could not measure depth ($d(t)=0$) and would need to assume a constant contact area ($A(t)=A_{\text{sensor}}$). We calculate contact stress $\sigma$ and object strain $\epsilon$, following Eq. 1 and Eq. 2. This model assumes symmetry between parallel fingers. We define $t=0$ to be the moment of first contact, where $F(t=0)$ must be greater than 0.75 N.
$\sigma(t)=\frac{F(t)}{A(t)}$ (1)
$\epsilon(t)=\frac{\Delta w(t)+2\Delta d(t)}{w(0)+2d(0)}=\frac{(w(t)+2d(t))-(w(0)+2d(0))}{w(0)+2d(0)}$ (2)
@@
Young’s modulus is defined as the ratio between elastic stress and strain [36]. Using measurements across grasping time $t$, we can fit a linear estimate for Young’s modulus $\hat{E}_{\text{elastic}}$ following Eq. 3.
$\bm{\sigma}=\hat{E}_{\text{elastic}}\bm{\epsilon}$ (3)
@@
Despite its simplicity, this method does not accurately calculate object stress or strain, especially if the object is larger than the sensor surface. Furthermore, it overlooks precise information about the deformation of contact geometry that is apparent through high-resolution tactile data.
@@
Figure 3: Parallel grasping and Hertzian mechanics. (a) Sphere of radius $R$ is grasped with soft GelSight sensors. Normal force $F$ is applied at width $w$. Tactile penetration depth $d$ and contact radius $a$ are measured with tactile sensors. Timestep $t=0$ is the moment of first contact over the grasp. (b) Normal force $F$ and maximum tactile depth $d$ measurements are collected across grasps for spheres of each material type and fit to $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
@@
Hertzian Contact: Hertzian contact theory models the deformation of elastic bodies in contact with each other. In a simple case, contact is assumed to be without friction or adhesion [10, 9]. Unlike previous methods, Hertzian mechanics closely analyzes geometry through contact. It has poreviously been applied to model soft contact in robotics with camera-based tactile sensors [23, 25, 22, 13].
@@
GelSight sensors are made with elastic silicone rubber [16]. Unless a grasped object is rigid, we expect both bodies to deform through contact. This implies the maximum measured tactile depth $d$ will be less than the relative displacement between the finger and object $d<\frac{\Delta w}{2}$, as depicted in Fig. 3a. An aggregate modulus $E^{*}$ describes the bodies’ combined resistance to deformation in Eq. 4, which depends on the Poisson’s ratio $\nu$ of each body.
@@
Hertzian solutions have been derived for a wide range of contact geometries [37]. We modeled the sensor as a half-space, which may be a reasonable approximation in most cases but could be less accurate for large, rigid objects that penetrate deeper into the sensor’s finite thickness. We modeled all grasped objects as spherical. This assumption is reasonable for convex objects, but may produce bounded errors over our geometrically diverse dataset. We expect that out neural network can learn to compensate for this intrinsic and repeatable geometric modeling error.
Since we assume contact profiles are spherical, we calculate contact radius $a$ based on observed total contact area $A$. Additionally, we calculate the apparent radius of the unknown object $R$ using contact area and gripper displacement [37]. Using these parameters, we can relate tactile depth $d$ to normal force $F$ as formulated in Eq. 6. This relationship is derived via the method of dimensionality reduction (MDR) in Appendix -A [9, 37, 38]. With equations established and measurements over the duration of a grasp, we can apply least-squares to retrieve a best fit value $\hat{E}^{*}$.
$a(t)=\sqrt{\frac{A(t)}{\pi}},\quad R(t)=\frac{a^{2}(t)}{\Delta w(t)}$ (5)
$\bm{d}=\frac{1-\nu_{\text{sensor}}^{2}}{E_{\text{sensor}}}{\left(\frac{3{{\hat{E}}^{*2}}\bm{F}}{32\bm{R}^{2}}\right)}^{\frac{1}{3}}\bm{a}$ (6)
With calculated $\hat{E}^{*}$, we can extract an estimated modulus for an unknown object $\hat{E}_{\text{hertz}}$ through Eq. 7, given known properties of our sensor $E_{\text{sensor}}=0.275$ MPa, $\nu_{\text{sensor}}=0.48$. We assume constant $\nu_{\text{obj}}=0.4$ based on the Poisson’s ratio of common materials [39]. After estimates are computed, a linear scaling is added to account for potential constant errors across the dataset.